OpenStudy (anonymous):

Is there a rule for the number of photons emitted when a matter particle and a corresponding anti-matter particle annihilate?

OpenStudy (experimentx):

hv = e = mc^2 hv = mc^2

OpenStudy (anonymous):

when a matter particle and a corresponding anti-matter particle annihilate they change completely into energy.. E=mc^2... this energy comes out as photons (say of frequency f),then nhf=E.. where n is no of photons..thus we can calculate the total no of photons..

OpenStudy (experimentx):

seems like i forgot the basics ... lol

OpenStudy (anonymous):

happens to everyone, someday or the other :)

OpenStudy (experimentx):

;)